OpenStudy (anonymous):

Given two points A and B, there is only one segment that you can name: AB. With three collinear points A, B, and C, there are three different segments that you can name: AB, AC, and BC. With five collinear points A, B, C, D, and E, how many different segments can you name?

OpenStudy (anonymous):

Using the network formula, we receive that: \[ \#P=\frac{n^2-n}{2} \]So, we plug in values for each necessary \(n\).

OpenStudy (anonymous):

One example, we have 3 points: \[ \#L=\frac{(3)^2-3}{2}=\frac{6}{2}=3 \]